OpenStudy (the_puppet1):

Substitute the given value of x inn the equation 3x + 7 = 1, and simplify the left sides of these equations. Answers in comments

OpenStudy (freckles):

what given value of x?

OpenStudy (freckles):

3x+7 means 3 times x then add 7 so just replace x with whatever x they gave you...

OpenStudy (anonymous):

Yea what is it or do u have to find out what x is?

OpenStudy (elise_a18):

x= -7 over 3

OpenStudy (the_puppet1):

I'm not done

OpenStudy (elise_a18):

oops actually -2

OpenStudy (anonymous):

Yep it is -2 without plugging anything in

OpenStudy (the_puppet1):

Here it is

OpenStudy (anonymous):

3x+7=1 3(-2)+7=1 -6+7=1 1=1